OpenStudy (anonymous):

Find the slope (3/8, -1) (-1/4, -1/4)

OpenStudy (anonymous):

slope m = rise/run

OpenStudy (shamil98):

slope = y2-y1/x2-x1

OpenStudy (anonymous):

\[Y _{2}-Y _{1}/ X _{2}-X_{1}\]

OpenStudy (anonymous):

so if im doing it correctly is it -1? @iambatman

OpenStudy (anonymous):

No your result is wrong

OpenStudy (anonymous):

Plug in the points with the slope formula we gave you

OpenStudy (anonymous):

y = -6/8 = -3/4 x = 3/4 ? @iambatman

OpenStudy (anonymous):

\[\frac{ -1/4-(-1) }{ -1/4-3/8 }\]

OpenStudy (anonymous):

what does that =

OpenStudy (anonymous):

3/4 4/8? @iambatman

OpenStudy (anonymous):

ok so divide them and what does it =?

OpenStudy (anonymous):

formula is y2-y1/ x2-x1

OpenStudy (anonymous):